0.3(w+70)=0.7(w+40)+25
Answer by checkley71(8403)

.3(W+70)=.7(W+40)+25 OR .3W+21=.7W+28+25 OR .3W+21=.7W+53 OR 21-53=.7W-.3W OR
-32=.4W OR -32/4=W OR -80=W
PROOF .3(-80+70)=.7(-80+40)+25 OR .3(-10)=.7(-40)+25 OR -3=-28+25 OR -3=-3